OpenStudy (anonymous):

You have a coupon for your favorite clothing store for $25 off any purchase of more than $50. The store is also running a 20%-off sale on its entire inventory. Let x be the original price, f(x) be the price with the $25 coupon applied, and g(x) be the price with the 20% discount applied. a. Write an expression for f(x). b. Write an expression for g(x). c. What would the expression (f o g)(x) represent? d. What would the expression (g o f)(x) represent? e. If the store allows you to apply both the 20% discount and the $25-off coupon, does it matter which you apply first? How do y

OpenStudy (anonymous):

***There are 5 parts.. :/ Can u please explain? Thank you :)

jimthompson5910 (jim_thompson5910):

If something costs x dollars and you apply a $25 coupon, how much is it now?

OpenStudy (anonymous):

x-25 ??

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

so f(x) = x - 25

OpenStudy (anonymous):

oh that simple? so that's my answer to part a? f(x)= x-25 ??

jimthompson5910 (jim_thompson5910):

If something costs x dollars, and you apply a 25% discount, then the final price will become x - 0.25x = 0.75x So g(x) = 0.75x

OpenStudy (anonymous):

ohhh i see... :) so now i apply that to part c??

jimthompson5910 (jim_thompson5910):

(f o g)(x) = f(g(x))

OpenStudy (anonymous):

so for part c i do this? (f o g)(x)=f(g(x)) ??? did i do that right?

OpenStudy (anonymous):

oh wait lol nvm u just wrote that.. my bad! haha so i do this now? f(g(x))=f(0.75x) ??

jimthompson5910 (jim_thompson5910):

good, keep going

OpenStudy (anonymous):

okay, so do i substitute f(x) now?

jimthompson5910 (jim_thompson5910):

f(x) = x - 25 f(0.75x) = 0.75x - 25 is that what you meant?

OpenStudy (anonymous):

yeah :) was that right?

jimthompson5910 (jim_thompson5910):

now do the reverse for part d

OpenStudy (anonymous):

wait so was this my answer for part c? (f o g)(x)= 0.75x - 25 ??

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

okay :) so for part d i get this? (g o f)(x)= g(f(x)) =g(x - 25) and I'm confused about the next step... did i do the previous steps correctly?

jimthompson5910 (jim_thompson5910):

you're doing great so far

OpenStudy (anonymous):

okay, so i have this?? g(x-25)= -25 +0.75 x ??? idk i'm confused :(

jimthompson5910 (jim_thompson5910):

g(x) = 0.75x g(x-25) = 0.75(x - 25) g(x-25) = 0.75x - 0.75*25 g(x-25) = 0.75x - 18.75 So g(f(x)) = 0.75x - 18.75

OpenStudy (anonymous):

ohhh that makes sense now! so i write out part d like this then?? (g o f)(x)= 0.75x - 18.75 ?? is that how i write the answer to part d?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

awesome! so i'm especially confused about part e... like how do i solve that? :O

jimthompson5910 (jim_thompson5910):

So part c was 0.75x - 25 part d was 0.75x - 18.75 which one offers the better discount?

OpenStudy (anonymous):

ohhh i use parts c and d to answer part e?

OpenStudy (anonymous):

is part c the 20% discount? or is part d? and which one is the $25 dollar off coupon?

jimthompson5910 (jim_thompson5910):

yes you compare parts c and d and see which one gives you the better discount

OpenStudy (anonymous):

oh in that case, i think the $25 dollar discount is better right?? part c is the better discount?? since it is -25, which is more than -18.75?? am i right?

OpenStudy (anonymous):

and am i right when i say that part c is the $25 off coupon and part d is the 20% off discount?

jimthompson5910 (jim_thompson5910):

parts c and d are a mix of both

OpenStudy (anonymous):

ohhh then how do i know which to apply first then? i'm lost now :(

jimthompson5910 (jim_thompson5910):

(f o g)(x) means apply g first, then apply f (g o f)(x) means apply f first, then apply g

OpenStudy (anonymous):

it seems to me that part c is the best since you save 6.25 more... i think... ohhh so would it matter or not?

jimthompson5910 (jim_thompson5910):

part c you are saving more, you got it

jimthompson5910 (jim_thompson5910):

so applying the 25% coupon first, then taking off 25 dollars gives you the better discount

OpenStudy (anonymous):

because of what i said?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

okay so can i phrase part e like this? e. It does matter, because applying the 20% discount would save you 18.75 and then you could use the $25 off coupon to save $6.25 more than you would if you had used the $25 off coupon first and then used the 20% off discount. When your total is higher, you save more from the 20% off discount. In this case, by using the 20% discount first and then using the $25 off coupon, you would save $6.25 more than you would have if you had used them the other way around. is that okay?

OpenStudy (anonymous):

*and is that correct?? is that the correct answer for part e???? :/ :/

jimthompson5910 (jim_thompson5910):

ugh, sry i misread above and it's not 25%...it's 20%....idk why i didn't catch that til now

jimthompson5910 (jim_thompson5910):

something wasn't matching up and i know why now

OpenStudy (anonymous):

oh okay, so is my answer incorrect??

OpenStudy (anonymous):

the one i typed for my answer as e above, is it wrong then @jim_thompson5910 ? :O

jimthompson5910 (jim_thompson5910):

yeah unfortunately it is

jimthompson5910 (jim_thompson5910):

idk why I'm seeing this now, my bad

OpenStudy (anonymous):

aww :( but i answered based off of 20% discount and $25 off :/

OpenStudy (anonymous):

it's okay :)

OpenStudy (anonymous):

so is answer b incorrect also?? and answers b,c, and d are incorrect then right??

jimthompson5910 (jim_thompson5910):

Here's what everything should be a) f(x) = x - 25 ------------- b) g(x) = 0.8x ------------- c) (f o g)(x) = 0.8x - 25 ------------- d) (g o f)(x) = 0.8x - 20 ------------- part c) still gives you the better discount

OpenStudy (anonymous):

okay so could u pls show me how you got b, c, and d?? like the steps?? i need to show work :/ and i also do not understand how u got those answers :/

jimthompson5910 (jim_thompson5910):

If you apply a 20% discount, then you subtract off 20% of the original cost. If the original price was x, then 20% of x is 0.2x So the new price is x - 0.2x = 0.8x Therefore, g(x) = 0.8x

OpenStudy (anonymous):

oh i see,... but what about parts c and d?

jimthompson5910 (jim_thompson5910):

f(x) = x - 25 f(g(x)) = g(x) - 25 f(g(x)) = 0.8x - 25 (f o g)(x) = 0.8x - 25

jimthompson5910 (jim_thompson5910):

g(x) = 0.8x g(f(x)) = 0.8(f(x)) g(f(x)) = 0.8(x-25) g(f(x)) = 0.8x-0.8*25 g(f(x)) = 0.8x-20 (g o f)(x) = 0.8x-20

OpenStudy (anonymous):

oh okay thanks!! so for part e it matters which i apply first right?

jimthompson5910 (jim_thompson5910):

it does matter yes

OpenStudy (anonymous):

okay, well i still think if i apply the 20% off (part d right?) coupon first, that's better, right? because i would save $5 more that way if i got 20% off first and then got $25 off right?? am i right? i would save $5 more?

jimthompson5910 (jim_thompson5910):

no, part c) is where you apply the 20% coupon first

OpenStudy (anonymous):

ohh i see... my bad! so applying the $25 off coupon first and then using the 20% discount would save you $5 more than if you used the $25 off coupon first? am i right?

jimthompson5910 (jim_thompson5910):

no, applying the 20% discount first then the $25 discount will save you 5 more dollars

OpenStudy (anonymous):

ohhh would that be using part c is the best choice? is that it?

jimthompson5910 (jim_thompson5910):

yes part c) is the best option if you want to save the most money

OpenStudy (anonymous):

okay so is this a good way to write my answer for part e? (lemme write it out)

OpenStudy (anonymous):

e. It does matter, because applying the 20% discount first and then using the $25 off coupon would save you $5 more than if you were to use the $25 off coupon first and then use the 20% discount. When your original total is higher, you save more from a percentage discount than a dollar discount, therefore, it matters since you would save $5 more if you use the 20% discount before you apply the $25 off coupon.

OpenStudy (anonymous):

does that look alright to you??

jimthompson5910 (jim_thompson5910):

That looks correct the last bit "When your original total is higher, you save more from a percentage discount than a dollar discount, therefore, it matters since you would save $5 more if you use the 20% discount before you apply the $25 off coupon." isn't 100% true Basically, you save more with applying the percent discount first, then taking off $25 every time Note: obviously the total must be above a certain mark for the discounts to apply (ie you're not going to have a discount that deducts the total price to $0 or below)

OpenStudy (anonymous):

so just take that part out?? and this will be my answer? e. It does matter, because applying the 20% discount first and then using the $25 off coupon would save you $5 more than if you were to use the $25 off coupon first and then use the 20% discount. Therefore, it matters which you apply first, because if you use the 20% discount first and then use the $25 off coupon, you will save $5 more. ***i fixed it a bit... does that look ok now? :)

jimthompson5910 (jim_thompson5910):

looks much better

OpenStudy (anonymous):



OpenStudy (anonymous):

those are the final, correct answers??! :O

jimthompson5910 (jim_thompson5910):

looks perfect

OpenStudy (anonymous):

awesome!! THank you sooooo much!! :D

jimthompson5910 (jim_thompson5910):

you're welcome